OpenStudy (twizttiez):

What is the solution to the system of equations? y=2x+6 3x-y=6 ( ? , ? )

OpenStudy (twizttiez):

@DanJS

OpenStudy (danjs):

oh wait

OpenStudy (danjs):

same thing as last one, the first equation is solved for y already, use that Y= in the second equation

OpenStudy (danjs):

3x - (2x+6) = 6

OpenStudy (twizttiez):

I think it's (0,-6)

OpenStudy (twizttiez):

So i was right?

OpenStudy (danjs):

wait no

OpenStudy (danjs):

see when you distribute this happens 3x - (2x+6) = 6 3x - 2x - 6 = 6 x = 6 +6 x=12

OpenStudy (twizttiez):

Oh so i was wrong?

OpenStudy (danjs):

yes, when you substitute Y into the second equation, it is a quantity in parenthesis 2x -1(2x+6) 2x - 2x - 6 ....not plus 6

OpenStudy (twizttiez):

Ok so what does that give me?

OpenStudy (twizttiez):

I'm sorry but i gotta finish this quick

OpenStudy (danjs):

x=12 y = 2*12 + 6

OpenStudy (danjs):

y=30

OpenStudy (twizttiez):

Ok thanks i have only 3 more

OpenStudy (twizttiez):

Or actually two